The enrolment in a school this year is 552. 'This is an increase of 15% over last year's enrolment. How many were enrolled last year ?
Answer by algebra hello(55)

let last year enrolment=x,then according to question x+x*15%=552 x+x*15/100=552 x(1+15/100)=552 x(115/100)=552 x=552*100/115 x=480